# Berkovich Motives

### v1

@@ 12. Relation to v-stacks
Now we have a natural forgetful functor $(R,R^{+})\mapsto R$ from perfectoid Huber pairs $(R,R^{+})$ with chosen norm towards Banach-$\mathbb{F}_{p}$-algebras. This has both adjoints. The right adjoint takes $R$ to $(R,R^{\circ})$ while the left adjoint takes $R$ to $(R,\widetilde{\mathbb{F}_{p}+R^{\circ\circ}})$ where $\widetilde{\mathbb{F}_{p}+R^{\circ\circ}}\subset R^{\circ}$ denotes the integral closure of $\mathbb{F}_{p}+R^{\circ\circ}$.
@@
The functors $(R,R^{+})\mapsto R$ and $R\mapsto(R,\widetilde{\mathbb{F}_{p}+R^{\circ}\circ})$ induce adjoint continuous maps of sites, and hence induce adjoint morphisms of $\infty$-topoi
@@
Most naturally occuring v-stacks are partially proper. This includes analyticifications of algebraic varieties, $\mathrm{Div}^{1}$, $\mathrm{Bun}_{G}$, Banach–Colmez spaces, etc. However, quasicompact open subsets are virtually never partially proper. Such quasicompact open subsets play a key technical role in [Sch17] as they are responsible for the compact generation of many categories, and many arguments proceed by reduction to compact objects. In the Berkovich setting, compact generation is not true anymore, but the categories are usually still compactly assembled, i.e. (in the stable setting) dualizable. This means that arguments involving reduction to compact objects have to be replaced by arguments involving reduction to compact morphisms.
@@
whose essential image consists of all overconvergent etale sheaves, i.e. those whose geometric fibres at higher-rank geometirc points $\mathrm{Spa}(C,C^{+})$ agree with the stalks at their generic point $\mathrm{Spa}(C,\mathcal{O}_{C})$.

### v2

@@ 12. Relation to v-stacks
[…]
@@
[…]
@@
Many naturally occuring v-stacks are partially proper. This includes analyticifications of algebraic varieties, $\mathrm{Div}^{1}$, $\mathrm{Bun}_{G}$, Banach–Colmez spaces, etc. However, quasicompact open subsets are virtually never partially proper. Such quasicompact open subsets play a key technical role in [Sch17] as they are responsible for the compact generation of many categories, and many arguments proceed by reduction to compact objects. In the Berkovich setting, compact generation is not true anymore, but the categories are usually still compactly assembled, i.e. (in the stable setting) dualizable. This means that arguments involving reduction to compact objects have to be replaced by arguments involving reduction to compact morphisms.
@@
whose essential image consists of all overconvergent étale sheaves, i.e. those whose geometric fibres at higher-rank geometirc points $\mathrm{Spa}(C,C^{+})$ agree with the stalks at their generic point $\mathrm{Spa}(C,\mathcal{O}_{C})$.